a 60$ sweater is on sale for 25% off what is the answer


Answer by stanbon(75887)

a $60 sweater is on sale for 25% off what is the answer?
----------------------
You had to pay 75% of the original price.
---
Ans: 0.75*60 = (3/4)60 = $45
------------
Cheers,
Stan H.
-----------